OpenStudy (anonymous):

Nathan is 6 feet tall and casts a shadow 2.5 feet long. At the same time, a lamppost casts a shadow 10 feet long. Write and solve a proportion to find the height of the lamppost. Please help and explain.

OpenStudy (anonymous):



OpenStudy (anonymous):

you have many choices for ratios as the triangles are similar you could say \[\frac{6}{2.4}=\frac{x}{10}\] or \[\frac{x}{6}=\frac{10}{2.4}\] but the answer will be the same in any case, it will be \[x=\frac{6\times 10}{2.4}\]